# High-pressure characterization of Ag3AuTe2: Implications for strain-induced band tuning

### v1

@@ Abstract
Recent band structure calculations have suggested the potential for band tuning in a chiral semiconductor, Ag3AuTe2, to zero upon application of negative strain. In this study, we report on the synthesis of polycrystalline Ag3AuTe2 and investigate its transport, optical properties, and pressure compatibility. Transport measurements reveal the semiconducting behavior of Ag3AuTe2 with high resistivity and an activation energy $E_{a}$ of 0.2 eV. The optical band gap determined by diffuse reflectance measurements is about three times wider than the experimental $E_{a}$. Despite the difference, both experimental gaps fall within the range of predicted band gaps by our first-principles DFT calculations employing the PBE and mBJ methods. Furthermore, our DFT simulations predict a progressive narrowing of the band gap under compressive strain, with a full closure expected at a strain of -4% relative to the lattice parameter. To evaluate the feasibility of gap tunability at such substantial strain, the high-pressure behavior of Ag3AuTe2 was investigated by $in$ $situ$ high-pressure X-ray diffraction up to 47 GPa. Mechanical compression beyond 4% of the lattice constant resulted in a pressure-induced structural transformation, indicating the potential for substantial gap modulation under extreme compression conditions.
@@ I Introduction
Metal chalcogenides have emerged as a remarkable class of materials, exhibiting unique electronic properties that make them suitable for exploring topological phenomena. Petzite or Ag3AuTe2 is a precious metal containing mineral, discovered approximately two centuries ago by PetzFaizan et al. (2017). It belongs to the uytenbogaardtite mineral group along with Ag3AuSe2 and Ag3AuS2. While Ag3AuSe2 and Ag3AuTe2 adopt the cubic $I4_{1}32$ space group, Ag3AuS2 takes on the tetragonal $P4_{1}22$ structure.Barton et al. (1978) Frueh et al.Frueh (1959) elucidated the crystal structure of natural Ag3AuTe2 by single crystal X-ray diffraction.
High-temperature powder X-ray diffraction experiment have revealed phase transitions in Ag3AuTe2 at 210 and 319 C, with an intermediate phase eluding identification.Cabri (1965); Frueh (1959) These transitions were corroborated by differential thermal analysis.Tavernier et al. (1967); Smit et al. (1970) Below the melting point of Ag3AuTe2 around 735 C, a continuous solid solution exists between the highest forms of Ag3AuTe2 and Ag2Te down to 319 CCabri (1965). The initial synthesis of Ag3AuTe2 was reported by Thompson,Thompson (1948) providing limited details on temperature or purity, except for conducting reactions in evacuated silica tubes. Young et al.Young et al. (2000) investigated the thermoelectric properties of synthetic Ag3AuTe2  noting low thermal conductivity and substantial electrical resistivity.
Recent computational predictions have pointed to a narrow and direct gap at $\Gamma$ Faizan et al. (2016); Young et al. (2000); Sánchez-Martínez et al. (2019), opening avenues for non-trivial topological behavior such as band crossing and inversion via band gap tuning. Sánchez-Martínez et al.Sánchez-Martínez et al. (2019) explored potential applications of Ag3AuSe2 and Ag3AuTe2 in dark matter detection, highlighting their tunable band gap through strain. Subsequent band closure of Ag3AuTe2 was predicted with a 4% reduction in the lattice parameter,Sánchez-Martínez et al. (2019) attainable through approaches such as chemical substitution and application of hydrostatic pressure application. Our preceding computational examination of the band gap tunability of Ag3AuSe2 demonstrated a narrowing of the gap under tensile strainWon et al. (2022), aligning with findings by Sánchez-Martínez et al.Sánchez-Martínez et al. (2019) However, such heightened strain on a material is subject to deformation, decomposition, or fracture. Therefore, a thorough assessment of the mechanical compressibility of Ag3AuTe2 is necessary to ascertain the feasibility of achieving a 4% compressive strain.
Previous studies on Ag3AuTe2Faizan et al. (2016); Young et al. (2000); Sánchez-Martínez et al. (2019) highlight a small direct gap with potential band closure under compressive strain, yet experimental band gap determination and the compressibility of Ag3AuTe2 remains scant. Here, we present the synthesis, transport, optical properties and mechanical compressibility of Ag3AuTe2 and explore the DFT-simulated band structure at varying lattice constants.
@@ II Methods
Powders of Ag (99.9% metals basis), Au (99.98% metals basis), and Te (99.8% metals basis) were loaded and sealed in an evacuated quartz ampule in a 3.05:1:1.9 ratio. Subsequently, the ampule underwent annealing in a muffle furnace at 770 C for a day, followed by rapid cooling to 200 C and a cooldown to room temperature over three days. This procedure yielded a brittle ingot exhibiting a metallic luster, covered partially with gold flakes on the surface. High-resolution powder X-ray diffraction (XRD) data were collected using the beamline 11-BM at the Advanced Photon Source (APS), Argonne National Laboratory (ANL).Lee et al. (2008) Data were collected at 288 K, using a wavelength of 0.45897 Å and the capillary geometry. Rietveld refinement was performed utilizing GSAS-II.Toby and Von Dreele (2013) Additionally, cube-shaped crystals of Ag3AuTe2 were picked out to confirm handedness and uniform chirality by single crystal X-ray diffraction (SCXRD). Transmission single crystal X-ray diffraction was measured in the Ag3AuTe2crystal using a Bruker D8 Venture diffractometer with Mo-K $alpha$ radiation at room temperature. Absorption correction was performed. The final refinements were done using SHELX.
High-pressure XRD measurements were performed at the 16-BMD-HPCAT beamline of the APS. For these experiments, finely ground sample material was loaded with a few grains of ruby and a Ne pressure-transmitting medium in a symmetric diamond anvil cell (DAC) with 300-$\mu$m culet diamonds and a Re gasket. The gaskets were indented to 60 $\mu$m, and a circular hole of 130 $\mu$m diameter was laser drilled to create the sample chamber. The Ne gas was loaded at the GSECARS facility at APS. The calibration of the sample center was carried out with a CeO2 standard, and the high-pressure data were collected up to 47 GPa with an incident beam of wavelength 0.41326 Å. Pressure at the sample region was measured and monitored by the shift of ruby fluorescence linesH.K.Mao and Bell (1986). The XRD images were collected with a Pilatus detector and integrated using Dioptas software Prescher and Prakapenka (2015).
Electrical resistivity measurements were performed using the 4-point contact probe method using a Quantum Design Physical Property Measurement System (PPMS). The Ag3AuTe2 ingot was cut and polished into a rod shape with dimensions 3.97 × 1.04 × 0.93 mm3 to facilitate transport measurements. Gold leads of diameter around 0.05 mm were attached to the sample with a two-part silver epoxy for the measurements.
@@ III.1 Growth and structure
Figure 2: Rietveld-refined synchrotron powder XRD pattern of Ag3AuTe2. The contributions from minor phases are shown in the figure and quantified in the text.
The room temperature crystal structure of petzite ($a=10.38$, $Z=8$) is analogous to fischesserite, Ag3AuSe2, as described in detail in our previous work.Won et al. (2022) The crystal structure of Ag3AuTe2, featuring a chiral $4_{1}$ Ag chain down the $\langle 100\rangle$ axis, is illustrated in Figure 1. Rietveld refinement to the Ag3AuTe2 PXRD pattern indicates the majority phase Ag3AuTe2 and minor impurities, as demonstrated in Figure 2. Phase fractions of Ag3AuTe2, Ag2Te, Ag doped Au2Te and Au are refined to 0.926, 0.035, 0.034, and 0.005 by weight, respectively, with R = 2.84%.
An equilibrated 3-component system can accomodate up to three phases in an isotherm, which denotes an disequilibrium in our final product. The Au-Ag2Te phase diagramCabri (1965) and the prior investigations into the thermal properties of Ag3AuTe2Tavernier et al. (1967); Smit et al. (1970); Cabri (1965); Frueh (1959) identify petzite as a line compound undergoing two phase transformations below its melting point. Despite our efforts to synthesize Ag3AuTe2 via a slow cool method from the melt, incorporating adjustments in cooling rates and elemental ratios, all attempts resulted in a mixture of Ag3AuTe2, Ag2Te, Au2Te, and Au, contradicting the reported phase diagram. There are certain ambiguities in the phase diagram proposed by Cabri et al. Cabri (1965), such as the dashed line above the melting point of Ag3AuTe2 labeled with a question mark. In addition, it is unclear whether the ”liquid” above the melting point is a uniform phase or molten impurities. Furthermore, the Ag3AuTe2 phase behavior near 200 C is ambiguous, with an intermediate phase mentioned but not indexed. This prompts questions about potential phase separation or deviation from the nominal 312 stoichiometry, leading to phase separation around the transition temperature. Consequently, the reformation of the lowest form of Ag3AuTe2 upon cooling may be challenging due to the difficulty in reaching an equilibrium at such low temperatures.
@@ III.2 High-Pressure X-ray Diffraction
Figure 3: (Top) $In$ $situ$ high-pressure XRD patterns collected at various pressures for Ag3AuTe2 ($\lambda=0.41326$). Le Bail fit for the orthorhombic ($Pnma$) structure at 47 GPa is displayed. (Bottom) Volume per formula unit of refined cells for Ag3AuTe2 plotted as a function of applied pressure. The multi-phase region around 16 GPa does not provide well-converged values for lattice parameters of the individual phases. The error bars for the volume are smaller than the symbol size.
Figure 4: Rietveld-refined lattice parameters and the corresponding $\sigma$ of compressed Ag3AuTe2 plotted against the applied pressure. Nearly -4% of strain is achieved at approximately 10 GPa.
@@
The high-pressure behavior of petzite (Ag3AuTe2) was investigated through in situ high-pressure XRD experiments. Figure 3 shows the X-ray diffraction patterns of Ag3AuTe2 collected at various representative pressures, reaching up to 47 GPa. The cubic structure of Ag3AuTe2 at ambient pressure remained stable up to 13 GPa, evidenced by the rightward shift of the most intense (110) peak 8.5. Beyond 13 GPa, the intensity of this prominent peak is decreased, indicating a pressure-induced phase change. Notably, changes above 10 were observed, with some peaks merging and new peaks emerging around 17.5. Rietveld refinement of lattice constants of Ag3AuTe2 demonstrated a gradual decrease with increasing pressure, reaching a maximum strain exceeding 4% at 13 GPa, as depicted in Figure 4. Rietveld analysis of the XRD patterns confirmed the retention of the ambient cubic phase of Ag3AuTe2 up to 13 GPa, with excellent fitting (wR = 2.60%). Even though the high-pressure XRD patterns collected between 13 GPa and 20 GPa show a mixture of both low-pressure and high-pressure phases of Ag3AuTe2, the significant changes observed in the diffraction peaks around 16 GPa imply structural transformation. Analysis commenced using the 47 GPa pattern for the high-pressure phase corresponds mostly to the high-pressure phase.
@@
We considered possible structural models from the pattern indexing and refinements using GSAS-II, LHPM Rietica, and QualXToby and Von Dreele (2013); Hunter (2000); Altomare et al. (2015) software programs as well as previously published papers for related compoundsBerri et al. (2013). Here, we consider an Al3CePt2-type hexagonal cell, $Pnma$ structure of CePd3In2,Nesterenko et al. (2004) and two other high-pressure polymorphs of Ag2Te: $C2/m$ from Zhu, et al.Zhu et al. (2016) and $Pnma$ from Zhang, et al.Zhang et al. (2015) Attempts to fit the observed diffraction patterns with the above candidate structures favored the Ag2Te-type structure, replicating most observed peaks at 47 GPa, as shown in Figure 3. The resulting cell parameters were a = 13.452(2) Å, b = 3.531(8) Å, c = 4.188(1) Å. Fitting lower-pressure patterns above 20 GPa with a similar cell revealed a volume collapse of around 6% around the phase transition. Therefore, it is likely that Ag3AuTe2 adopts the Pnma-type symmetry at high pressure. Ab initio enthalpy calculations on the proposed structural models would further shed light on their stability, especially since the number of observed diffraction peaks is small.
@@
The bulk modulus of the low pressure cubic and high pressure orthorhombic phases was obtained by fitting the $P$-$V$ data to a third-order Vinet equation of state Vinet et al. (1987), as shown in Figure 3. The fits provided a bulk modulus value of $B_{0}$ = 52(7) GPa with a derivative $B_{0}^{\prime}$ = 9.2 for the low-pressure phase and $B_{0}$ = 68(1) GPa with $B_{0}^{\prime}$ = 4.8 for the high-pressure phase. The compressibility of Ag3AuTe2 is comparable to other Ag ternary chalcogenides reported in the literature such as AgSbTe2 and AgBiSe2. Berri et al. (2013); Ko et al. (2014); Elafy et al. (2019) Understanding the mechanical compressibility of Ag3AuTe2 is crucial in light of the previous computational findings suggesting the possibility of band engineering and crossing under compressive strain. Further investigation into the high-pressure band gap dynamics of Ag3AuTe2 using complementary techniques such as $in~{}situ$ high-pressure UV-Vis and resistivity measurements may provide valuable insights into its topological nature.
@@ III.3 Electronic properties
Figure 5: Electrical resistivity of Ag3AuTe2 as a function of temperature indicating a semiconductor-like temperature dependence. The inset Arrhenius plot was obtained by applying a linear regression model with the least-squares method, which returned an activation energy of approximately 0.20 eV.
@@
Figure 5 shows the measured electrical resistivity of Ag3AuTe2 as a function of temperature, accompanied by the corresponding inset Arrhenius plot. The temperature dependence of electron transport indicates a semiconductor behavior with high resistivity, nearing the megaohm range at around 100 K. The magnitude of the Arrhenius activation energy aligns well with the previously proposed DFT-calculated band gaps.Faizan et al. (2016); Sánchez-Martínez et al. (2019) Young et al.Young et al. (2000) reported similar transport measurement results, albeit with a slightly larger experimental $E_{a}$ of 0.3 eV and no evidence for sample purity. Differences in the estimated activation energy may be attributed to imperfections in sample geometry, inhomogeneity, and the presence of possible in-gap defects in our sample. Additionally, a resistivity kink near 275 K in Figure 5, also observed by Young et al.Young et al. (2000) with no explanation, is likely an artifact from the freezing of adsorbed water.
Transport measurements confirm that Ag3AuTe2 exhibits a semiconductor behavior with a band gap of at least 0.2 eV, accounting for potential defect hopping effects. Accurate band gap determination is essential for evaluating petzite’s potential for strain-induced crossings of the valence and conduction bands. Subsequent DFT analysis explores the band structure and potential band gap modulation through lattice parameter adjustments.
Electronic properties were computationally investigated by density functional theory (DFT) band structure calculations. Figure 6 illustrates the calculated band structure of Ag3AuTe2 with a reduction of lattice constant up to 4% compressive strain, using the Perdew-Burke-Ernzerhof parameterization (PBE) and modified Becke-Johnson (mBJ), including spin-orbit coupling (SOC). Calculations without SOC are provided in the supplemental material. SOC breaks Kramer’s degeneracy and splits the band structure for different spin components, but it does not induce qualitative differences in the band dispersion and the gap. Table 1 summarizes the results, indicating that at ambient pressure ($a=10.38$), the indirect band gap between the valence band near $N$ and the conduction band around $\Gamma$ is 0.093 and 0.603 eV for PBE and mBJ functionals, respectively. mBJ calculations predict that negative strain (compression) decreases the band gap while conserving the overall shape and features of the band structure, and the material remains to be an insulator at the maximal compressive strain (-4%). The PBE functional, known to underestimate band gaps, also predicts decreasing band gaps with negative strain. However, the conduction and valence band overlap near $\Gamma$ along the $\Gamma-H$ cut before reaching -4% strain, transitioning the material into a metal under pressure. The SOC splitting strengthens with increased strain and the predicted SOC effect by mBJ functional is less significant than that using PBE method. Our PBE results are mostly consistent with previous DFT studiesSánchez-Martínez et al. (2019); Gosálbez-Martínez et al. (2023). However, we expect the mBJ functional to provide more accurate calculation of the band structure in general, and the shape and splitting of the conduction band around $\Gamma$, in particular, are qualitatively different from the PBE predictions.

### v2

@@ Abstract
Recent band structure calculations have suggested the potential for band tuning in a chiral semiconductor, Ag3AuTe2, to zero upon application of negative strain. In this study, we report on the synthesis of polycrystalline Ag3AuTe2 and investigate its transport, optical properties, and pressure compatibility. Transport measurements reveal the semiconducting behavior of Ag3AuTe2 with high resistivity and an activation energy $E_{a}$ of 0.2 eV. The optical band gap determined by diffuse reflectance measurements is about three times wider than the experimental $E_{a}$. Despite the difference, both experimental gaps fall within the range of predicted band gaps by our first-principles DFT calculations employing the PBE and mBJ methods. Furthermore, our DFT simulations predict a progressive narrowing of the band gap under compressive strain, with a full closure expected at a strain of -4% relative to the lattice parameter. To evaluate the feasibility of gap tunability at such substantial strain, the high-pressure behavior of Ag3AuTe2 was investigated by in situ high-pressure X-ray diffraction up to 47 GPa. Mechanical compression beyond 4% resulted in a pressure-induced structural transformation, indicating the possibilities of substantial gap modulation under extreme compression conditions.
@@ I Introduction
Metal chalcogenides have emerged as a remarkable class of materials, exhibiting unique electronic properties that make them suitable for exploring topological phenomena. Petzite or Ag3AuTe2 is a mineral, discovered by Petz about two centuries agoFaizan et al. (2017). It belongs to the uytenbogaardtite mineral group along with Ag3AuSe2 and Ag3AuS2. While Ag3AuSe2 and Ag3AuTe2 adopt the cubic $I4_{1}32$ space group, Ag3AuS2 takes on the tetragonal $P4_{1}22$ structure.Barton et al. (1978) Frueh et al.Frueh (1959) elucidated the crystal structure of natural Ag3AuTe2 by single crystal X-ray diffraction (XRD).
High-temperature powder XRD experiment revealed phase transitions in Ag3AuTe2 at 210 and 319 C, with an intermediate phase eluding identification.Cabri (1965); Frueh (1959) These transitions were corroborated by differential thermal analysis.Tavernier et al. (1967); Smit et al. (1970) The initial synthesis of Ag3AuTe2 was reported by Thompson,Thompson (1948) providing limited details on temperature or purity, except for conducting reactions in evacuated silica tubes. Young et al.Young et al. (2000) investigated the thermoelectric properties of synthetic Ag3AuTe2  noting low thermal conductivity and large electrical resistivity.
Recent computational predictions have pointed to a narrow and direct gap at $\Gamma$ Faizan et al. (2016); Young et al. (2000); Sánchez-Martínez et al. (2019), opening possibilities for non-trivial topological behavior such as band crossing and inversion via band gap tuning. Sánchez-Martínez et al.Sánchez-Martínez et al. (2019) explored potential applications of Ag3AuSe2 and Ag3AuTe2 in dark matter detection, highlighting their tunable band gap through strain. Subsequent band closure of Ag3AuTe2 was predicted with a 4% reduction in the lattice parameter,Sánchez-Martínez et al. (2019) attainable through approaches such as chemical substitution and application of hydrostatic pressure. Our preceding computational examination of the band gap tunability of Ag3AuSe2 demonstrated a narrowing of the gap under tensile strainWon et al. (2022), aligning with findings by Sánchez-Martínez et al.Sánchez-Martínez et al. (2019) However, such heightened strain on a material is subject to deformation, decomposition, or fracture. Therefore, a thorough assessment of the mechanical compressibility of Ag3AuTe2 is necessary to evaluate the feasibility of band engineering by achieving -4% strain.
Previous studies on Ag3AuTe2Faizan et al. (2016); Young et al. (2000); Sánchez-Martínez et al. (2019) highlight a small direct gap with potential band closure under compressive strain, yet experimental determination of the band gap and compressibility remains scant. Here, we report the synthesis, transport, optical properties and mechanical compressibility of Ag3AuTe2 and explore the DFT-simulated band structure at varying lattice constants.
@@ II Methods
Powders of Ag (99.9% metals basis), Au (99.98% metals basis), and Te (99.8% metals basis) were loaded and sealed in an evacuated silica ampule in a 3.05:1:1.9 ratio. Subsequently, the ampule underwent annealing in a muffle furnace at 770 C for a day, followed by rapid cooling to 200 C and a cool-down to room temperature over three days. This procedure yielded a brittle ingot exhibiting a metallic luster and covered partially with gold flakes on the surface. High-resolution powder X-ray diffraction (PXRD) data were collected using the beamline 11-BM at the Advanced Photon Source (APS), Argonne National Laboratory (ANL).Lee et al. (2008) Data were collected at 288 K, using a wavelength of 0.45897 Å and the capillary geometry. Rietveld refinement was performed utilizing GSAS-II.Toby and Von Dreele (2013)
High-pressure XRD measurements were performed at the 16-BMD-HPCAT beamline of the APS. Finely ground sample was loaded with a few grains of ruby and a Ne pressure-transmitting medium in a symmetric diamond anvil cell (DAC) with 300-$\mu$m culet diamonds and a Re gasket. The gaskets were indented to 60 $\mu$m, and a circular hole of 130 $\mu$m diameter was laser drilled to create the sample chamber. The Ne gas was loaded at the GSECARS facility at APS. The calibration of the sample center was carried out with a CeO2 standard, and the high-pressure data were collected up to 47 GPa with an incident beam of wavelength 0.41326 Å. Pressure at the sample region was measured and monitored by the shift of ruby fluorescence linesH.K.Mao and Bell (1986). The XRD images were collected with a Pilatus detector and integrated using Dioptas software Prescher and Prakapenka (2015).
Electrical resistivity measurements were performed by the 4-point contact probe method using the Quantum Design Physical Property Measurement System (PPMS). The Ag3AuTe2 ingot was cut and polished into a rod with dimensions 3.97 × 1.04 × 0.93 mm3 to facilitate transport measurements. Gold leads were attached to the sample with a two-part silver epoxy for the measurements.
@@ III.1 Growth and structure
Figure 2: Rietveld-refined synchrotron PXRD pattern of Ag3AuTe2. The contributions from minor phases are shown in the figure and quantified in the text.
The room temperature crystal structure of petzite ($a=10.38$, $Z=8$) is analogous to fischesserite, Ag3AuSe2, as described in detail in our previous work.Won et al. (2022) The crystal structure of Ag3AuTe2 features a chiral $4_{1}$ Ag chain down the $\langle 100\rangle$ axis, as illustrated in Figure 1. Our Ag3AuTe2 sample was characterized by PXRD and analysed using Rietveld method. As shown in Figure 2, Bragg peaks were well-indexed by the cubic petzite structure and some minor impurities (wR = 2.84%). Phase fractions of Ag3AuTe2, Ag2Te, Ag doped Au2Te and Au were refined to 0.926, 0.035, 0.034, and 0.005 by weight, respectively.
An equilibrated 3-component system can accommodate up to three phases in an isotherm, which denotes an disequilibrium in our final product. The Au-Ag2Te phase diagramCabri (1965) and previous studies on the thermal properties of Ag3AuTe2Tavernier et al. (1967); Smit et al. (1970); Cabri (1965); Frueh (1959) identify petzite as a line compound undergoing two phase transformations below its melting point. Despite our efforts to synthesize Ag3AuTe2 via a slow cooling method from the melt, incorporating adjustments in cooling rates and elemental ratios, all attempts resulted in a mixture of Ag3AuTe2, Ag2Te, Au2Te, and Au, contradicting the reported phase diagram. There are certain ambiguities in the phase diagram proposed by Cabri et al. Cabri (1965), such as the dashed line above the melting point of Ag3AuTe2 labeled with a question mark. In addition, it is unclear whether the ”liquid” above the melting point is a uniform phase or molten impurities. Furthermore, the Ag3AuTe2 phase behavior near 200 C is ambiguous, with an intermediate phase mentioned but not indexed. This prompts questions about potential phase separation or deviation from the nominal 312 stoichiometry, leading to phase separation around the transition temperature. Consequently, the reformation of the lowest form of Ag3AuTe2 upon cooling may be challenging due to the difficulty in reaching an equilibrium at such low temperatures.
Figure 3: (Top) in situ high-pressure XRD patterns of Ag3AuTe2 collected at various pressures ($\lambda=0.41326$). Le Bail fit for the orthorhombic ($Pnma$) structure at 47 GPa is displayed. (Bottom) Volume per formula unit of refined cells for Ag3AuTe2 plotted as a function of applied pressure. The multi-phase region around 16 GPa does not provide well-converged values for lattice parameters of the individual phases. The error bars for the volume are smaller than the symbol size.
Figure 4: Rietveld-refined lattice parameters and the corresponding $\sigma$ of compressed Ag3AuTe2 are plotted against the applied pressure. Nearly -4% of strain is achieved at 10 GPa.
@@ III.2 High-Pressure X-ray Diffraction
The high-pressure behavior of Ag3AuTe2 was investigated through in situ high-pressure XRD experiments. Figure 3 depicts the XRD patterns collected at various representative pressures, reaching up to 47 GPa. The cubic structure of Ag3AuTe2 at ambient pressure remained stable up to 13 GPa, demonstrated by the rightward shift of the most intense (110) peak at 8.5. Beyond 13 GPa, this prominent peak drops in intensity, indicating a pressure-induced phase change. Notably, a number of peaks merge around 10 and new peaks emerge near 17.5. Rietveld analysis confirmed the retention of the ambient cubic phase up to 13 GPa and the resulting refined lattice constant $a$ showed a gradual decrease with increasing pressure, reaching a maximum strain exceeding -4% at 13 GPa, as shown Figure 4. Even though the patterns collected between 13 and 20 GPa show a mixture of both low- and high-pressure phases of Ag3AuTe2, the significant changes observed in the diffraction peaks around 16 GPa imply structural transformation. Analysis commenced using the 47 GPa pattern for the high-pressure phase corresponds mostly to the high-pressure phase.
@@
We considered possible structural models from the pattern indexing and refinements using GSAS-II, LHPM Rietica, and QualXToby and Von Dreele (2013); Hunter (2000); Altomare et al. (2015) software programs as well as previously published papers for related compoundsBerri et al. (2013). Here, we consider an Al3CePt2-type hexagonal cell, $Pnma$ structure of CePd3In2,Nesterenko et al. (2004) and two high-pressure polymorphs of Ag2Te: $C2/m$ from Zhu, et al.Zhu et al. (2016) and $Pnma$ from Zhang, et al.Zhang et al. (2015) Attempts to fit the observed diffraction patterns with the above candidate structures favored the Ag2Te-type structure, replicating most observed peaks at 47 GPa, as shown in Figure 3. The resulting cell parameters were a = 13.452(2) Å, b = 3.531(8) Å, c = 4.188(1) Å. Fitting the patterns above 20 GPa with a similar cell resulted in a volume collapse of approximately 6% around the phase transition, adding weight to the likeliness that Ag3AuTe2 adopts the Pnma-type symmetry at high pressure. Ab initio enthalpy calculations on the proposed structural models would further shed light on their stability, especially since the number of observed diffraction peaks is small.
@@
The bulk modulus of the low-pressure (cubic) and high-pressure (orthorhombic) phases was obtained by fitting the $P$-$V$ data to a third-order Vinet equation of state Vinet et al. (1987), as shown in Figure 3. The fits provided a bulk modulus value of $B_{0}$ = 52(7) GPa with a derivative $B_{0}^{\prime}$ = 9.2 for the cubic phase and $B_{0}$ = 68(1) GPa with $B_{0}^{\prime}$ = 4.8 for the orthorhombic phase. The compressibility of Ag3AuTe2 is comparable to other reported Ag ternary chalcogenides such as AgSbTe2 and AgBiSe2.Berri et al. (2013); Ko et al. (2014); Elafy et al. (2019) The high strain tolerance of Ag3AuTe2 bestow prospects for further investigation into the high-pressure band gap dynamics in light of the previous computational findings which suggested the possibility of strain-induced band engineering. Complementary techniques such as in situ high-pressure UV-Vis and resistivity measurements would provide valuable insights into its topological nature.
@@ III.3 Electronic properties
Figure 5: Electrical resistivity of Ag3AuTe2 is plotted against temperature. The inset Arrhenius plot was obtained by applying a linear regression model with the least-squares method, which returned an activation energy of approximately 0.20 eV.
Electronic properties of Ag3AuTe2 were examined experimentally by transport measurements. Figure 5 shows the electrical resistivity of Ag3AuTe2 as a function of temperature with the corresponding inset Arrhenius plot generated using the activation law. The material exhibits a semiconductor behavior and high resistivity, nearing the megaohm range at around 100 K. The magnitude of the activation energy $E_{a}$ aligns well with the previously proposed DFT-calculated band gaps.Faizan et al. (2016); Sánchez-Martínez et al. (2019) Young et al.Young et al. (2000) reported similar experimental results, albeit with a slightly larger $E_{a}$ of 0.3 eV and no evidence for sample purity. Differences in the estimated $E_{a}$ are within the typical error range for resistivity measurements stemming from factors such as imperfections in sample geometry, inhomogeneity, and possible in-gap defects. Additionally, a resistivity kink near 275 K in Figure 5, also observed by Young et al.Young et al. (2000) with no explanation, is likely an artifact from the freezing of adsorbed water.
Transport measurements confirm that Ag3AuTe2 shows semiconductor-like temperature dependence and a band gap of at least 0.2 eV, accounting for potential defect hopping effects. To evaluate petzite’s potential for strain-induced band crossing, accurate determination as well as prediction of band gap is important. Subsequent DFT analysis explores the band structure and possible band gap modulation through lattice parameter adjustments.
@@
Electronic properties were computationally investigated by density functional theory (DFT) band structure calculations. Figure 6 illustrates the calculated band structure of Ag3AuTe2 with a reduction of lattice constant up to 4% compressive strain, using the Perdew-Burke-Ernzerhof parameterization (PBE) and modified Becke-Johnson (mBJ), including spin-orbit coupling (SOC). Calculations without SOC are provided in the supplementary material. SOC breaks Kramer’s degeneracy and splits the band structure for different spin components, but it does not induce qualitative differences in the band dispersion and the gap. Table 1 summarizes the results, indicating that at ambient pressure ($a=10.38$), the indirect band gap between the valence band near $N$ and the conduction band around $\Gamma$ is 0.093 and 0.603 eV for PBE and mBJ functionals, respectively. mBJ calculations predict that negative strain (compression) decreases the band gap while conserving the overall shape and features of the band structure, and the material remains to be an insulator at the maximal compressive strain (-4%). The PBE functional, known to underestimate band gaps, also predicts decreasing band gaps with negative strain. However, the conduction and valence band overlap near $\Gamma$ along the $\Gamma-H$ cut before reaching -4% strain, transitioning the material into a metal under pressure. The SOC splitting strengthens with increased strain and the predicted SOC effect by mBJ functional is less significant than that using PBE method. Our PBE results are mostly consistent with previous DFT studiesSánchez-Martínez et al. (2019); Gosálbez-Martínez et al. (2023). However, we expect the mBJ functional to provide more accurate calculation of the band structure in general, and the shape and splitting of the conduction band around $\Gamma$, in particular, are qualitatively different from the PBE predictions.